OpenStudy (anonymous):

If the point P(-3/5, y)lies on the unit circle and P is the third quadrant, what does y equal?

OpenStudy (anonymous):

@misty1212 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

P lies on the unit circle, so we know that x^2 + y^2 = 1

jimthompson5910 (jim_thompson5910):

x = -3/5 is given to you use this to find y

jimthompson5910 (jim_thompson5910):

P is the quadrant 3, so y will be negative

jimthompson5910 (jim_thompson5910):

does all that make sense? or no?

OpenStudy (anonymous):

so far yes

jimthompson5910 (jim_thompson5910):

so you need to solve for y \[\Large x^2 + y^2 = 1\] \[\Large \left(\frac{-3}{5}\right)^2 + y^2 = 1\] \[\Large \frac{9}{25} + y^2 = 1\] \[\Large y = ???\]

OpenStudy (anonymous):

I got 4/5 and -4/5

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi!

OpenStudy (misty1212):

what is the most famous right triangle of all time?

OpenStudy (anonymous):

45 45 90?

OpenStudy (misty1212):

lol those are angles silly i meant the one with the nice integer sides

jimthompson5910 (jim_thompson5910):

y = -4/5 is your answer because remember P is in quadrant 3

jimthompson5910 (jim_thompson5910):

y is negative in quadrant 3

OpenStudy (misty1212):



OpenStudy (misty1212):

divide everything by \(5\) and you get this 

OpenStudy (misty1212):

since your first coordinate was \(\frac{3}{5}\) the second one had to be \(\frac{4}{5}\) or maybe \(-\frac{4}{5}\) depending on the quadrant